OpenStudy (anonymous):

Years- 2000| 2001| 2002| 2003| 2004| 2005| 2006| 2007| 2008| 2009 Population- 1605| 1728| 1776| 1903| 2072| 2050| 2141| 2150| 2318| 2363 a. Make a scatter plot of the data, draw a line of best fit, and then find the equation of the line of best fit. Show and explain your work. b. Describe what the slope of the line of best fit represents. c. Use the equation to predict the population in the year 2013. Show your work.

OpenStudy (anonymous):

I know I need to find the slope so I choose two years with their populations and calculate... m= 2000- 2001/ 1605- 1728= 1/ 123 1/123= slope I know how to get that far I just don't know where to go from there. Can someone help me figure this out??

OpenStudy (anonymous):

Use excel: col 1: years col 2: population. You can plot the data as a scatter plot using the plot tool

OpenStudy (anonymous):

just chosing two years and taking the differences in population will give you a limited view and may not reflect the entire data

OpenStudy (anonymous):

Look at all the slopes and get the average rate of change for the population. The assumption is that it will continue to be this average rate so use that rate to calculate the subsequent years

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem, let me know if you have other questions!